there are 90 students , 60 play clarinet and 40 play the trombone, 10 only play the trombone. find the probability of a student playing the trombone or the clarinet.
Answer by ewatrrr(24785)


Hi
there are 90 students , 60 play clarinet and 40 play the trombone,
10 only play the trombone: that is 30 play both.
P(trombone or clarinet) = 40/90 + 60/90 - 30/90 = 70/90 = 7/9
P(A or B) = P(A) + P(B) - P(A and B)